# Causal Structure Learning in Hawkes Processes with Complex Latent Confounder Networks

### v1

@@ Appendix E Accounting for Self-Looped Observed Subprocesses under Latent Confounder Influence
For any subprocess $O_{1}$ that is not a single-node subprocess isolated from all others and itself, we prove the equivalence of the four statements step by step.
(1) $\Leftrightarrow$ (2): If $\mathcal{P}_{\mathcal{G}}$ is the parent cause set of $O_{1}$ in the summary graph, by construction of the window causal graph, it equivalent to that the corresponding lagged variable set $\mathbf{P}_{v}$ contains all direct parent variables of $O_{1}^{(n)}$. This follows from the fact that, in the window graph, directed edges exist from the effective lag variables of each parent subprocess to $O_{1}^{(n)}$. Moreover, by definition of the parent cause set, $\mathcal{P}_{\mathcal{G}}$ is minimal with this property.
(2) $\Leftrightarrow$ (3): If $\mathbf{P}_{v}$ contains all direct parents of $O_{1}^{(n)}$ in the window graph, by the Markov property of DAGs, $\mathbf{P}_{v}$ d-separates $O_{1}^{(n)}$ from all other observed variables in $\mathbf{O}_{v}\backslash\left(\mathbf{P}_{v}\cup\{O_{1}^{(n)}\}\right)$. Reversly, if $\mathbf{P}_{v}$ d-separates $O_{1}^{(n)}$ from all other observed variables in $\mathbf{O}_{v}\backslash\left(\mathbf{P}_{v}\cup\{O_{1}^{(n)}\}\right)$, by the Granger causality-events in the future cannot causally influence events in the past, $\mathbf{P}_{v}$ should contain all direct parents of $O_{1}^{(n)}$ in the window graph. Moreover, by definition of the parent cause set, $\mathcal{P}_{\mathcal{G}}$ is minimal with this property.
(3) $\Leftrightarrow$ (4): By applying Lemma˜3.2, the d-separation between $O_{1}^{(n)}$ and the rest of the variables, conditioned on $\mathbf{P}_{v}$, is equivalent to the rank constraint:
@@
(4) $\Leftrightarrow$ (1): Assume the rank condition holds for $\mathbf{P}_{v}$. By Lemma˜3.2, this implies that $\mathbf{P}_{v}$ d-separates $O_{1}^{(n)}$ from all other variables in the window graph. Translating back to the summary graph, this implies that $\mathcal{P}_{\mathcal{G}}$ is the minimal parent cause set of $O_{1}$, as no smaller set can block all paths to $O_{1}$.
@@
Appendix J Preliminaries for Proofs of Propositions˜3.5, 3.7 and 3.8
To establish this result, we rely on the concepts of trek separation (t-separation) and d-separation introduced by [50], which provide powerful tools for analyzing latent structures in linear causal models.
Definition J.1 (Trek [50]).
@@
Theorem J.3 (Trek separation for directed graphical models [50]).
@@
Corollary J.4 (T-separation and D-separation [50]).
@@
Appendix K Proof of Proposition˜3.5
@@
Suppose there exists a latent confounder $L_{1}$ that is one common parent cause in the parent cause set of $\{O_{1},O_{2}\}$, and that $L_{1}$ together with $\{O_{1},O_{2}\}$ makes them locally independent of other subprocesses.
Given that $L_{1}$ and its paths to $O_{1}$ and $O_{2}$ satisfy Definition˜3.4, the contribution of $L_{1}$ to both $O_{1}$ and $O_{2}$ in the window graph occurs through the same number of latent intermediates, resulting in an aligned contribution across time lags. In this setup, the influence of $L_{1}$ will appear as a shared component across the observed variables $\{O_{i}^{(j)}\}^{i\in\{1,2\}}_{j\in\{n-m,\dots,n\}}$.
Consider the window graph with $m$ considered effective lag variables. Following the logic of trek separation, in the window graph with $m$ effective lag variables, the minimal choke set $\mathbf{C_{A}}$ that t-separates $O_{1}^{(n)},O_{2}^{(n)}$} from the rest is given by:
@@
Thus, by Theorem˜J.3, the generic rank of the cross-covariance matrix is bounded above by $|\mathbf{C_{A}}|=2m+m=3m$, where $2m$ comes from observed lag variables of $\{O_{1},O_{2}\}$ and $m$ comes from latent lag variables of $L_{1}$. However, due to the structure of the excitation function $\phi_{ij}(s)=a_{ij}w(s)$, the latent subprocess $L_{1}$ contributes effectively as a single shared component across all its lag variables, reducing the effective rank from $m$ to $1$.
@@
By construction of the window graph (Eq. 2), if there were no latent confounder between $O_{1}$ and $O_{2}$, the rank would be at most $2m$, corresponding to the observed lag variables of $O_{1}$ and $O_{2}$. The observed rank being strictly $2m+1$ thus implies the presence of an additional latent variable influencing both $O_{1}$ and $O_{2}$.
Due to the rank faithfulness assumption (˜2), such a rank elevation uniquely corresponds to a latent subprocess $L_{1}$ acting as a parent cause of both $O_{1}$ and $O_{2}$. Furthermore, for the rank increment to be exactly one, the causal paths from $L_{1}$ to $O_{1}$ and $O_{2}$ must satisfy the symmetric path situation (Definition˜3.4): i.e., the paths only involve intermediate latent subprocesses of the same depth without self-loops, ensuring that the contribution of $L_{1}$ introduces a single additional rank component shared by both $O_{1}$ and $O_{2}$ at the same temporal lag level.
@@
Appendix L Proof of Theorem˜3.7
@@
($\Leftarrow$) If such a parent cause set $\mathcal{P}^{\prime}_{\mathcal{G}}$ exists, the rank condition holds.
@@
  • $\mathcal{P}^{\prime}_{\mathcal{G}}$ is a subset of the parent cause set of $N_{1}$.
@@
  • For each $L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}$, the relationship between $L_{i}$ and its observed effects $\{\mathcal{D}e(N_{1}),\mathcal{D}e(L_{i})\}$ satisfies Definition˜3.4.
@@
Given the symmetric path structure (Definition˜3.4), each latent confounder $L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}$ contributes exactly one shared latent component, as the influence propagates through symmetric, acyclic paths. Due to the specific excitation function $\phi_{ij}(s)=a_{ij}w(s)$, this results in precisely one rank contribution per latent subprocess, regardless of the number of lagged variables.
@@
If no latent confounders existed, or if $\mathcal{P}^{\prime}_{\mathcal{G}}$ were not part of the parent cause set of $N_{1}$, the rank would be exactly $|\mathbf{O}_{v1}|$, solely contributed by the lagged variables of observed surrogates and both the current and lagged variables of their siblings.
Since all possible observed surrogates of $N_{i}$ in $\mathcal{O}_{\mathcal{G}}$ have been identified, the extra deficiency of rank (i.e., $|\mathbf{O}_{v2}|$) thus directly implies the existence of latent subprocesses contributing shared rank-one components. By the rank faithfulness, this observed rank pattern is only consistent with the existence of latent subprocesses $\{L_{i}\}_{L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}}$ that act as confounders between $\mathcal{D}e(N_{1})$ and their respective observed effects, and these latent subprocesses are members of the parent cause set of $N_{1}$.
For the rank deficit per latent subprocess to be exactly one, the contribution from each latent subprocess must propagate through symmetric acyclic paths, consistent with Definition˜3.4, ensuring a single rank-one component contribution per latent subprocess. Moreover, the inclusion of the observed surrogates and their siblings ensures that no alternative paths can explain the dependency patterns. Thus, $\mathcal{P}^{\prime}_{\mathcal{G}}$ must be the subset of parent causes, satisfying the conditional local independence of $N_{1}$ given $\mathcal{S}_{\mathcal{G}}$.
@@
Appendix M Proof of Theorem˜3.8
@@
  • $L_{1}$ and its observed effects $\{\mathcal{D}e(N_{1}),\mathcal{D}e(N_{2})\}$ satisfy Definition˜3.4.
By the Definition˜3.4, the causal influence from $L_{1}$ to $\{\mathcal{D}e(N_{1}),\mathcal{D}e(N_{2})\}$ is symmetric and only propagates through the same number of intermediate latent subprocesses without self-loops. Under this condition, the contributions of $L_{1}$ to the observed surrogates $\{\mathcal{D}e(N_{1}),\mathcal{D}e(N_{2})\}$ appear as a rank-one component across the lagged variables of these subprocesses, aligned in time.
@@
Formally, following the arguments for Proposition˜3.5, the rank of $\Sigma_{\mathbf{A}_{v},\mathbf{B}_{v}}$ is determined by the minimal set of choke points that t-separate $\mathbf{A}_{v}$ from $\mathbf{B}_{v}$ in the window graph. Given the assumed structure:
@@
  • The influence from $L_{1}$ propagates symmetrically to both $\mathcal{D}e(N_{1})$ and $\mathcal{D}e(N_{2})$ through acyclic paths composed exclusively of latent subprocesses, per Definition˜3.4. As a result, due to the excitation function $\phi_{ij}(s)=a_{ij}w(s)$, the total rank contribution from $L_{1}$ is exactly one.
@@
In the absence of any latent confounder, the maximum possible rank would be $|\mathbf{O}_{v1}|$, corresponding solely to the contributions of the lagged variables of $\{\mathcal{D}e(N_{1}),\mathcal{D}e(N_{2})\}$ and both the current and lagged variables of their observed siblings. The observed rank being exactly $|\mathbf{O}_{v1}|+1=|\mathbf{A}_{v}|-1$ implies the existence of an additional latent source influencing both $N_{1},N_{2}$} and their observed surrogates.
Due to the rank faithfulness, this increment must be attributed to a unique latent subprocess $L_{1}$ that acts as a confounder for $N_{1}$ and $N_{2}$. Moreover, the fact that the rank increment is only one implies that the paths from $L_{1}$ to $N_{1},N_{2}$} must satisfy the symmetric and acyclic conditions in Definition˜3.4, ensuring that the influence of $L_{1}$ is captured as a rank-one shared component at the observed surrogates level.
@@
Appendix N Proof of Theorem˜4.1
@@
We prove the theorem by considering the two cases separately: (i) the system contains no latent subprocesses, and (ii) the system contains latent subprocesses that satisfy Definition˜3.4.
@@
In this case, the system consists solely of observed subprocesses $\mathcal{O}_{\mathcal{G}}$. Since there are no latent confounders, Phase I alone is sufficient for identifiability. This follows directly from Proposition˜3.3, which ensures that for each observed subprocess, its parent cause set can be uniquely identified by checking the rank condition of the relevant cross-covariance matrices. Specifically, since all subprocesses are observed and no latent subprocesses confound their relationships, the rank condition provides a unique solution. Thus, the entire causal graph can be identified solely through Phase I.
Case (ii): Presence of latent subprocesses satisfying Definition˜3.4.
@@
  • Phase I iteratively identifies the parent cause set for each subprocess (including both observed and previously discovered latent subprocesses) whose parent cause set is fully contained in the current set of known subprocesses. By Proposition˜3.3 and Theorem˜3.7, this identification is guaranteed when no latent confounders intervene or when latent confounders are already represented by their observed surrogates.
  • Phase II handles the discovery of new latent confounder subprocesses by systematically applying Propositions˜3.5 and 3.8. The identifiability is guaranteed under the assumption that all latent confounders and their associated observed effects satisfy Definition˜3.4. This condition ensures that each latent confounder contributes a unique, identifiable rank-1 pattern in the cross-covariance matrix of its observed surrogates and their siblings, enabling its detection through the rank conditions established in the theorems.
@@
The algorithm alternates between Phase I and Phase II. Since each iteration either identifies a new parent cause set or discovers a new latent subprocess, and given the finite number of subprocesses (including latent ones), the algorithm must eventually terminate.
@@
  • All observed subprocesses will eventually have their parent cause sets identified through Phase I.
  • All latent subprocesses satisfying Definition˜3.4 will be identified through Phase II and incorporated into the active set for further investigation.
  • The recursive application of the identification theorems ensures that no causal relationships (either between observed, latent, or between observed and latent) will remain unidentified under the assumptions.
Thus, under the given assumptions, the entire causal graph consisting of both observed subprocesses and latent confounders can be identified. This completes the proof. ∎
@@
6:    if  $(\mathcal{A}_{\mathcal{G}}\cup\mathcal{O}_{\mathcal{G}},\mathcal{P}^{\prime}_{\mathcal{G}},N_{1})$ satisfies Propositions˜3.3 and 3.7 then
@@
4:  if  $(\mathcal{A}_{\mathcal{G}}\cup\mathcal{O}_{\mathcal{G}},\mathcal{Y}_{\mathcal{G}})$ satisfies Propositions˜3.5 and 3.8 then
@@
For each component process $N_{1}\in\mathcal{A}_{\mathcal{G}}$, we evaluate subsets of $\mathcal{T}_{\mathcal{G}}$ starting from subsets of size 1 up to the size of $\mathcal{T}_{\mathcal{G}}$, stopping when the test result is positive. In the worst case, for each $N_{1}$, we need to evaluate all subsets of $\mathcal{T}_{\mathcal{G}}$, which requires $\sum_{k=1}^{m}\binom{m}{k}$ tests. For one subprocess $N_{1}\in\mathcal{A}_{\mathcal{G}}$, if its parent cause set is found, $\mathcal{A}_{\mathcal{G}}$ is updated. After that, the algorithm will restart to go over all the subprocesses in $\mathcal{A}_{\mathcal{G}}$ to make sure no parent cause set of subprocesses in $\mathcal{A}_{\mathcal{G}}$ can be found. In the worst case, the algorithm find parent cause set for the last component process in $\mathcal{A}_{\mathcal{G}}$ each time. The complexity of Phase I is upper bounded by: $\mathcal{O}\left(n!\sum_{k=1}^{m}\binom{m}{k}\right)$.

### v2

@@ Appendix E Accounting for Self-Looped Observed Subprocesses under Latent Confounder Influence
For any subprocess $O_{1}$, we prove the equivalence of the four statements step by step.
(1) $\Leftrightarrow$ (2): If $\mathcal{P}_{\mathcal{G}}$ is the parent-cause set of $O_{1}$ in the summary graph, by construction of the window causal graph, it equivalent to that the corresponding lagged variable set $\mathbf{P}_{v}$ contains all direct parent variables of $O_{1}^{(n)}$. This follows from the fact that, in the window graph, directed edges exist from the effective lagged variables of each parent subprocess to $O_{1}^{(n)}$. Moreover, by definition of the parent-cause set, $\mathcal{P}_{\mathcal{G}}$ is minimal with this property.
(2) $\Leftrightarrow$ (3): If $\mathbf{P}_{v}$ contains all parent variables of $O_{1}^{(n)}$ in the window graph, by the Markov property of DAGs, $\mathbf{P}_{v}$ d-separates $O_{1}^{(n)}$ from all other observed variables in $\mathbf{O}_{v}\backslash\left(\mathbf{P}_{v}\cup\{O_{1}^{(n)}\}\right)$. Reversly, if $\mathbf{P}_{v}$ d-separates $O_{1}^{(n)}$ from all other observed variables in $\mathbf{O}_{v}\backslash\left(\mathbf{P}_{v}\cup\{O_{1}^{(n)}\}\right)$, by the Granger causality-events in the future cannot causally influence events in the past, $\mathbf{P}_{v}$ should contain all parent variables of $O_{1}^{(n)}$ in the window graph. Moreover, by definition of the parent-cause set, $\mathcal{P}_{\mathcal{G}}$ is minimal with this property.
(3) $\Leftrightarrow$ (4): By applying Lemma˜4.2, the d-separation between $O_{1}^{(n)}$ and the rest of the variables, conditioned on $\mathbf{P}_{v}$, is equivalent to the rank constraint:
@@
(4) $\Leftrightarrow$ (1): Assume the rank condition holds for $\mathbf{P}_{v}$. By Lemma˜4.2, this implies that $\mathbf{P}_{v}$ d-separates $O_{1}^{(n)}$ from all other observed variables in the window graph. Translating back to the summary graph, this implies that $\mathcal{P}_{\mathcal{G}}$ is the minimal parent-cause set of $O_{1}$, as no smaller set can block all paths to $O_{1}$.
@@
Appendix J Preliminaries for Proofs of Propositions˜4.5, 4.7 and 4.8
To establish this result, we rely on the concepts of trek separation (t-separation) and d-separation introduced by [Sullivant et al., 2010], which provide powerful tools for analyzing latent structures in linear causal models.
Definition J.1 (Trek [Sullivant et al., 2010]).
@@
Theorem J.3 (Trek separation for directed graphical models [Sullivant et al., 2010]).
@@
Corollary J.4 (T-separation and D-separation [Sullivant et al., 2010]).
@@
Appendix K Proof of Proposition˜4.5
@@
Suppose there exists a latent confounder $L_{1}$ that is one common parent cause in the parent-cause set of $\{O_{1},O_{2}\}$, and that $L_{1}$ together with $\{O_{1},O_{2}\}$ makes them locally independent of other subprocesses.
Given that $L_{1}$ and its paths to $O_{1}$ and $O_{2}$ satisfy Definition˜4.4, the contribution of $L_{1}$ to both $O_{1}$ and $O_{2}$ in the window graph occurs through the same number of latent intermediates, resulting in an aligned contribution across time lags. In this setup, the influence of $L_{1}$ will appear as a shared component across the observed variables $\{O_{i}^{(j)}\}^{i\in\{1,2\}}_{j\in\{n-m,\dots,n\}}$.
Consider the window graph with $m$ considered effective lags. Following the logic of trek separation, in the window graph with $m$ effective lagged variables, the minimal choke set $\mathbf{C_{A}}$ that t-separates $O_{1}^{(n)},O_{2}^{(n)}$} from the rest is given by:
@@
Thus, by Theorem˜J.3, the generic rank of the cross-covariance matrix is bounded above by $|\mathbf{C_{A}}|=2m+m=3m$, where $2m$ comes from observed lagged variables of $\{O_{1},O_{2}\}$ and $m$ comes from latent lagged variables of $L_{1}$. However, due to the structure of the excitation function $\phi_{ij}(s)=a_{ij}w(s)$, the latent subprocess $L_{1}$ contributes effectively as a single shared component across all its lagged variables, reducing the effective rank from $m$ to $1$.
@@
By construction of the window graph (Eq. 2), if there were no latent confounder between $O_{1}$ and $O_{2}$, the rank would be at most $2m$, corresponding to the observed lagged variables of $O_{1}$ and $O_{2}$. The observed rank being strictly $2m+1$ thus implies the presence of an additional latent variable influencing both $O_{1}$ and $O_{2}$.
Due to the rank faithfulness assumption (˜2), such a rank elevation uniquely corresponds to a latent subprocess $L_{1}$ acting as a parent cause of both $O_{1}$ and $O_{2}$. Furthermore, for the rank increment to be exactly one, the causal paths from $L_{1}$ to $O_{1}$ and $O_{2}$ must satisfy the symmetric path situation (Definition˜4.4): i.e., the paths only involve intermediate latent subprocesses of the same depth without self-loops, ensuring that the contribution of $L_{1}$ introduces a single additional rank component shared by both $O_{1}$ and $O_{2}$ at the same temporal lag level.
@@
Appendix L Proof of Theorem˜4.7
@@
($\Leftarrow$) If such a parent-cause set $\mathcal{P}^{\prime}_{\mathcal{G}}$ exists, the rank condition holds.
@@
  • $\mathcal{P}^{\prime}_{\mathcal{G}}$ is a subset of the parent-cause set of $N_{1}$.
@@
  • For each $L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}$, the relationship between $L_{i}$ and its observed effects $\{\mathcal{D}e(N_{1}),\mathcal{D}e(L_{i})\}$ satisfies Definition˜4.4.
@@
Given the symmetric path structure (Definition˜4.4), each latent confounder $L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}$ contributes exactly one shared latent component, as the influence propagates through symmetric, acyclic paths. Due to the specific excitation function $\phi_{ij}(s)=a_{ij}w(s)$, this results in precisely one rank contribution per latent subprocess, regardless of the number of lagged variables.
@@
If no latent confounders existed, or if $\mathcal{P}^{\prime}_{\mathcal{G}}$ were not part of the parent-cause set of $N_{1}$, the rank would be exactly $|\mathbf{O}_{v1}|$, solely contributed by the lagged variables of observed surrogates and both the current and lagged variables of their siblings.
Since all possible observed surrogates of $N_{i}$ in $\mathcal{O}_{\mathcal{G}}$ have been identified, the extra deficiency of rank (i.e., $|\mathbf{O}_{v2}|$) thus directly implies the existence of latent subprocesses contributing shared rank-one components. By the rank faithfulness, this observed rank pattern is only consistent with the existence of latent subprocesses $\{L_{i}\}_{L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}}$ that act as confounders between $\mathcal{D}e(N_{1})$ and their respective observed effects, and these latent subprocesses are members of the parent-cause set of $N_{1}$.
For the rank deficit per latent subprocess to be exactly one, the contribution from each latent subprocess must propagate through symmetric acyclic paths, consistent with Definition˜4.4, ensuring a single rank-one component contribution per latent subprocess. Moreover, the inclusion of the observed surrogates and their siblings ensures that no alternative paths can explain the dependency patterns. Thus, $\mathcal{P}^{\prime}_{\mathcal{G}}$ must be the subset of parent causes, satisfying the conditional local independence of $N_{1}$ given $\mathcal{S}_{\mathcal{G}}$.
@@
Appendix M Proof of Theorem˜4.8
@@
  • $L_{1}$ and its observed effects $\{\mathcal{D}e(N_{1}),\mathcal{D}e(N_{2})\}$ satisfy Definition˜4.4.
By the Definition˜4.4, the causal influence from $L_{1}$ to $\{\mathcal{D}e(N_{1}),\mathcal{D}e(N_{2})\}$ is symmetric and only propagates through the same number of intermediate latent subprocesses without self-loops. Under this condition, the contributions of $L_{1}$ to the observed surrogates $\{\mathcal{D}e(N_{1}),\mathcal{D}e(N_{2})\}$ appear as a rank-one component across the lagged variables of these subprocesses, aligned in time.
@@
Formally, following the arguments for Proposition˜4.5, the rank of $\Sigma_{\mathbf{A}_{v},\mathbf{B}_{v}}$ is determined by the minimal set of choke points that t-separate $\mathbf{A}_{v}$ from $\mathbf{B}_{v}$ in the window graph. Given the assumed structure:
@@
  • The influence from $L_{1}$ propagates symmetrically to both $\mathcal{D}e(N_{1})$ and $\mathcal{D}e(N_{2})$ through acyclic paths composed exclusively of latent subprocesses, per Definition˜4.4. As a result, due to the excitation function $\phi_{ij}(s)=a_{ij}w(s)$, the total rank contribution from $L_{1}$ is exactly one.
@@
We know that parent-cause sets of all inferred latent confounder processes in $\mathcal{N}_{\mathcal{G}}$ remain unidentified even after applying Theorem˜4.7. In the absence of any new latent confounder, the maximum possible rank would be $|\mathbf{O}_{v1}|$, corresponding solely to the contributions of the lagged variables of $\{\mathcal{D}e(N_{1}),\mathcal{D}e(N_{2})\}$ and both the current and lagged variables of their observed siblings. The observed rank being exactly $|\mathbf{O}_{v1}|+1=|\mathbf{A}_{v}|-1$ implies the existence of an additional latent source influencing both $N_{1},N_{2}$} and their observed surrogates.
Due to the rank faithfulness, this increment must be attributed to a unique latent subprocess $L_{1}$ that acts as a confounder for $N_{1}$ and $N_{2}$. Moreover, the fact that the rank increment is only one implies that the paths from $L_{1}$ to $N_{1},N_{2}$} must satisfy the symmetric and acyclic conditions in Definition˜4.4, ensuring that the influence of $L_{1}$ is captured as a rank-one shared component at the observed surrogates level.
@@
Appendix N Proof of Theorem˜5.1
@@
We prove the theorem by considering the two cases separately: (i) the system contains no latent subprocesses, and (ii) the system contains latent subprocesses.
@@
In this case, the system consists solely of observed subprocesses $\mathcal{O}_{\mathcal{G}}$. Since there are no latent confounders, Phase I alone is sufficient for identifiability. This follows directly from Proposition˜4.3, which ensures that for each observed subprocess, its parent-cause set can be uniquely identified by checking the rank condition of the relevant cross-covariance matrices. Thus, the entire causal graph can be identified solely through Phase I.
Case (ii): Presence of latent subprocesses.
@@
  • Phase I iteratively identifies the parent-cause set for each subprocess (including both observed and previously discovered latent subprocesses) whose parent-cause set is fully contained in the current set of known subprocesses. By Proposition˜4.3 and Theorem˜4.7, this identification is guaranteed when no unknown latent confounders intervene or when latent confounders are already represented by their observed surrogates.
  • Phase II handles the discovery of new latent confounder subprocesses by systematically applying Propositions˜4.5 and 4.8. The identifiability is guaranteed under the condition that all latent confounder subprocesses and their associated observed effects satisfy Definition˜4.4. This condition uniquely ensures that each latent confounder subprocess contributes a unique, identifiable rank-1 pattern in the cross-covariance matrix of its observed surrogates and their siblings, enabling its detection through the rank conditions established in the theorems. The latent cofounder subprocesses that do not satisfy Definition˜4.4 remain unidentified.
@@
The algorithm alternates between Phase I and Phase II. Since each iteration either identifies a new parent-cause set or discovers a new latent subprocess, and given the finite number of subprocesses (including latent ones), the algorithm must eventually terminate.
@@
  • All observed subprocesses will eventually have their parent-cause sets identified through Phase I.
  • All latent subprocesses satisfying Definition˜4.4 will be identified through Phase II and incorporated into the active set for further investigation.
  • The recursive application of the identification theorems ensures that no causal relationships (either between observed, latent, or between observed and latent) will remain unidentified under the conditions.
  • If Definition˜4.4 fails for any latent, the algorithm terminates without fabricating that latent or any edges it would entail, thereby returning only the identifiable portion of the causal graph (sound abstention).
Thus, under excitation function $\phi_{ij}(s)=a_{ij}w(s)$ and rank faithfulness, the entire causal graph consisting of both observed subprocesses and latent confounders can be identified. This completes the proof. ∎
@@
6:     if $(\mathcal{A}_{\mathcal{G}}\cup\mathcal{O}_{\mathcal{G}},\mathcal{P}^{\prime}_{\mathcal{G}},N_{1})$ satisfies Propositions˜4.3 and 4.7 then
@@
4:   if $(\mathcal{A}_{\mathcal{G}}\cup\mathcal{O}_{\mathcal{G}},\mathcal{Y}_{\mathcal{G}})$ satisfies Propositions˜4.5 and 4.8 then
@@
For each component process $N_{1}\in\mathcal{A}_{\mathcal{G}}$, we evaluate subsets of $\mathcal{T}_{\mathcal{G}}$ starting from subsets of size 1 up to the size of $\mathcal{T}_{\mathcal{G}}$, stopping when the test result is positive. In the worst case, for each $N_{1}$, we need to evaluate all subsets of $\mathcal{T}_{\mathcal{G}}$, which requires $\sum_{k=1}^{m}\binom{m}{k}$ tests. For one subprocess $N_{1}\in\mathcal{A}_{\mathcal{G}}$, if its parent-cause set is found, $\mathcal{A}_{\mathcal{G}}$ is updated. After that, the algorithm will restart to go over all the subprocesses in $\mathcal{A}_{\mathcal{G}}$ to make sure no parent-cause set of subprocesses in $\mathcal{A}_{\mathcal{G}}$ can be found. In the worst case, the algorithm find parent-cause set for the last component process in $\mathcal{A}_{\mathcal{G}}$ each time. The complexity of Phase I is upper bounded by: $\mathcal{O}\left(n!\sum_{k=1}^{m}\binom{m}{k}\right)$.